# 22footnotetext: Corresponding author

### v1

@@ 3.1.2. Variants of U-Net
Ruan et al. (Ruan and Xiang, 2024) proposed VM-UNet where images are converted into tokens using a patch embedding block. The encoder of the network contains two VSS blocks with patch merging block and skip connection in each layer of the encoder. The decoder contains a patch expanding block followed by a VSS block with skip connections added from the encoder. Finally, the representations from the decoder are then passed on to the final projection layer which reconstructs the image back into its original size i.e., the number of classes. Figure 13 illustrates the UNet-based architecture of VM-UNet.
Figure 13. Architecture of VM-UNet (Ruan and Xiang, 2024)
Archit et al. (Archit and Pape, 2024) introduced ViM-UNet which uses ViM as its base Mamba. ViM-UNet features ViM encoder with bidirectional SSM layers, while its decoder copies the design of UNEt TRansformers (UNETR) (Hatamizadeh et al., 2021) using convolution and transposed convolution layers. To enhance efficiency, ViM-UNet introduces flexibility in its encoder sizes, such as tiny and small. In contrast, the traditional UNet uses a standard CNN for biomedical segmentation. UNet relies heavily on skip connections between corresponding encoder and decoder blocks to capture better features. However, ViM-UNet takes a different approach by excluding these skip connections and using ViM as the encoder. This key architectural change allows ViM-UNet to achieve efficient global feature extraction and improved segmentation performance, without the need for the skip connections that are important in the traditional UNet.
@@
Inspired by VM-UNet (Ruan and Xiang, 2024), Zhang et al. (Zhang et al., 2024d) proposed the second version called VM-UNet-V2. The model differs from VM-UNet by not using skip connections between subsequent encoder to decoder layers as shown in Figure 14. Instead, features from each block of encoder are fused using Semantics and Details Infusion block (SDI block). VSS captures contextual information within images, while SDI module improves the integration of both low-level and high-level features, leading to a comprehensive understanding of image features. By combining these elements, VM-UNetV2 maximizes the potential of SSMs within the UNet, thus offering a more efficient and powerful solution for segmentation tasks.
Figure 14. Architecture of VM-UNet V2(Zhang et al., 2024d)
@@
Wang et al. (Wang et al., 2024a) proposed LKM-UNet, a novel architecture designed for efficient 2D and 3D medical image segmentation. LKM-UNet uses strengths of Mamba to achieve superior performance in both local and global modeling with linear time complexity. The use of large windows within SSM improves the receptive field compared to CNNs and Transformers. The network architecture includes a hierarchical and bidirectional Mamba block, enhancing spatial modeling capabilities by integrating Pixel-level (PiM) and Patch-level (PaM) SSMs. LKM-UNet excels in capturing both fine-grained details and long-range dependencies from data. The model achieves efficient feature extraction and segmentation through a U-shaped network with encoder-decoder blocks, layers for downsampling and upsampling along with skip connections.
Figure 15. Architecture of H-vmunet (Wu et al., 2024a)
Wu et al. (Wu et al., 2024a) proposed High-order Vision Mamba UNet (H-vmunet) which enhances the 2D-selective-scan (SS2D) mechanism and also introduced higher-order interactions to reduce redundant information and improve extraction of local features. Traditional architecture like CNN and ViT face limitations in processing long sequences and capturing local features respectively. The proposed architecture features a U-shaped structure with six layers which incorporates High-order Visual State Space (H-VSS) module and it applies higher-order SS2D from layers three to six. The inclusion of Spatial Attention Bridge (SAB) and Channel Attention Bridge (CAB) modules in the architecture as shown in Figure 15 further enhances multilevel and multiscale information fusion which is crucial for capturing detailed medical image features.
3.1.3. Hybrid Architectures
@@
CNNs often face challenges in capturing long-range dependencies due to their inherent focus on local features and computational complexity. SSMs have the ability to handle long sequences of data. The combination of Mamba with convolution plays a huge role in capturing local spatial information along with capturing long range dependencies from medical images. Some of the papers that demonstrate these hybrid approaches are as follows: Ma et al. (Ma et al., 2024) introduced U-Mamba, a hybrid CNN-SSM architecture integrating Mamba blocks within the encoder of U-Net, demonstrating superior performance over traditional CNN-based and transformer-based segmentation networks across various modalities and segmentation targets. Wang et al. (Wang et al., 2024d) proposed Mamba-U-Net which integrates pure Vision mamba in the U-Net along with linear embedding and VSS block in the model. Xu et al. (Xu, 2024) developed HC-Mamba, which uses dilated convolution followed by depth-wise separable convolution along with Mamba. It improves the receptive field and reduces the parameters of the model. Another innovative model, SegMamba designed by Xing et al. (Xing et al., 2024), mimics a Transformer-based U-Net architecture, utilizing a series of TS-Mamba block and downsample block with residual connections inspired by He et al. (He et al., 2015). TS-Mamba employs gated spatial convolution to analyze spatial relationships between 3D features, followed by the Tri-oriented Spatial Mamba (TOM) block, which replaces the traditional multi-headed attention layers in transformers. Gong et al. (Gong et al., 2024) proposed nnMamba, which features the Mamba-In-Convolution with Channel-Spatial Siamese (MICCSS) block. The backbone architecture of nnMamba leverages the MICCSS block to maintain computational efficiency. This helps in improving the model’s representational capacity and improving the performance on tasks that require a deep understanding of visual data.
Figure 16. Architecture of Swin-UMamba (Liu et al., 2024b)
Liu et al. (Liu et al., 2024b) proposed Swin-UMamba which is a hybrid architecture that combines VSS with UNet for medical image segmentation tasks. Swin-UMamba detailed in Figure 16 integrates a Mamba-based encoder pretrained on ImageNet with a series of VSS blocks and upsample blocks. The architecture of Swin-UMamba includes patch embedding, VSS blocks, patch merging, up-sample blocks, residual blocks, and a 1x1 convolution for segmentation output which combines Mamba with traditional segmentation network components. By combining the strengths of Mamba-based models with segmentation network structures, Swin-UMamba achieves superior results in medical image segmentation tasks, highlighting its effectiveness in improving pretrained models for improved segmentation accuracy. Figure 16 details the architecture of Swin-UMamba.
@@
Wang et al. (Wang and Ma, 2024) proposed Weak-Mamba-UNet which uses a CNN-based UNet for extracting local features from data, a Swin Transformer-based SwinUNet for understanding global context and a VMamba based Mamba-UNet for modeling long-range dependencies. The results of Weak-Mamba-UNet shows that it solves inaccurate predictions caused by other approaches. Kirillov et al. proposed Segment Anything Model (SAM) (Kirillov et al., 2023) which is originally done for 2D images. Inspired by this, Wang et al. (Wang et al., 2024b) proposed Efficiently Adapting Segment Anything Model (SAM) which shows superior performance on 3D medical images. The authors have modified the encoder of SAM by adding Tri-Plane Mamba (TP-Mamba) block on top of each ViT and the decoder of the model uses 3D convolution with instance normalization (Ulyanov et al., 2017) and GELU activation function (Hendrycks and Gimpel, 2023). The authors have also used LoRA (Hu et al., 2022) to adapt weights of Multi-Head self Attention module. TP-Mamba shows superior performance compared to transformer based models, U-Mamba (Ma et al., 2024) and other adapter-based algorithms for SAM such as SA-Med (Zhang and Liu, 2023), MA-SAM (Chen et al., 2023). Figure 17 gives an overview of pretrained SAM with adapters on 3D segmentation, ViT blocks from pretrained SAM and the architecture of TP-Mamba.
Figure 17. Architecture of Tri-Plane Mamba (Wang et al., 2024b)
Zhang et al. (Zhang et al., 2024a) proposed HMT-UNet which combines MambaVision Mixer with the transformer’s self attention block for segmentation of medical images. MambaVision mixer is a modification of the Mamba layer, where the first casual convolution is replaced with regular convolution and passed on to SSM. Initially, the encoder of UNet consists of four layers where the initial two layers uses two $3\times 3$ convolution with stride of two and downsampling uses batch normalized $3\times 3$ convolution with stride of two which scales down the resolution of the spatial dimension by half and doubles the channel. The layer 3 and layer 4 contains MambaVision Mixer and transformer’s self attention block. In the decoder block, the spatial dimensions are doubled and channel size is halved. The final stages of decoder use the same configuration as initial stages of encoder but instead of downsampling the image as in encoder, it linearly upsamples the image.
@@
VMRNN proposed by Tang et al. (Tang et al., 2024a) is a recurrent cell that incorporates VSS block within an Long Short Term Memory (LSTM) (Hochreiter and Schmidhuber, 1997) network. Mamba’s integration into spatial-temporal forecasting based on vision allows robust sequential modeling. The paper introduces two novel architectures: VMRNN-B and VMRNN-D. These architectures excel at extracting spatiotemporal features, establishing a new strong foundation for spatiotemporal forecasting. Spatiotemporal predictive learning differs from traditional image-level vision tasks by predicting future video frames based on past sequences. VMRNN tackles this challenge by processing individual frames, segmenting them into smaller patches. These patches gets flattened before feeding them into a patch embedding layer for initial processing. The VMRNN layer then leverages these transformed patches alongside past states to capture the crucial spatiotemporal features necessary for predicting the next frame. The hybrid architecture of VSS and LSTM is attributed to the model’s capability to learn and leverage global spatial dependencies with linear complexity, enabling a more refined understanding of spatiotemporal dynamics.
@@
Ding et al. (Ding et al., 2024) combines Mamba with Graph Neural Network (GNN) to capture global and local tissue spatial relationships respectively in Whole Slide Images (WSI). The model leverages a message-passing GNN, specifically a Graph Attention Network (GAT) proposed by Velickovic et al. (Velickovic et al., 2017), to process a hierarchical graph constructed from both cell-level and tile-level graphs. Graph convolution operation is performed using GAT which assigns importance scores to neighboring nodes, allowing the model to focus on the most relevant information for the current node during the convolution process.
@@
Nasiri et al. (Nasiri-Sarvi et al., 2024) proposed Vim4Path which used ViM architecture within the DINO framework proposed by Caron et al. (Caron et al., 2021), for learning representations in computational pathology. ViM is modified to accept arbitrary input image sizes using positional embedding interpolation, making it adaptable within DINO for Self-Supervised Learning (SSL). The study benchmarks the Camelyon16 dataset (Bandi, [n. d.]), extracting image patches from Whole Slide Images (WSIs) without labels for training the ViM encoder using the DINO framework. Yang et al. (Yang et al., 2024d) introduced the MambaMIL framework, which combines the Mamba framework with Multiple Instance Learning (MIL) to enhance long sequence modeling in computational pathology. The core component of the MambaMIL framework is the Sequence Reordering Mamba (SR-Mamba), which is designed to be aware of the order and distribution of instances within long sequences. Mamba is used to enable each instance through a compressed hidden state to interact with previously scanned instances, facilitating effective modeling of long sequences while reducing computational complexity. The MambaMIL framework partitions tissue regions into a sequence of patches, maps these patches into instance features, reduces feature dimension through linear projection, utilizes stacked SR-Mamba modules for handling long sequences, and finally employs an aggregation module to obtain bag-level representations for downstream tasks.
3.2. Scanning
@@
Attention mechanisms, especially self-attention have a quadratic time complexity causing computational costs to grow quadratically with sequence length. In contrast, scanning operations generally have linear time complexity, making them more efficient for long sequences. The scan operation involves calculating an array, like the prefix sum, where each value is determined by using the previously calculated value and the current input. Similarly, the recurrent form of SSM can be viewed as a scan operation. Scanning is a crucial component in mamba, especially when handling multidimensional inputs. The selection of the scanning mechanism in Mamba models is crucial as it enhances efficiency and provides important information. LABEL:fig:scanning_part2 provides visualization of various scanning mechanisms employed in mamba-based architectures. This visualization highlights the diversity of scanning approaches integrated into mamba models. Table 1 summarizes the various scanning mechanisms and the models associated with each mechanism. The scanning methodologies used in mamba models are detailed as follows:
  1. (1)
@@
    Bidirectional Scan : In bidirectional scan (forward and backward scan) (Zhu et al., 2024), after tokenizing image patches, they are processed through the forward SSM. Simultaneously, the same tokenized representations of images are independently processed through the backward SSM. This scanning mechanism primarily used in ViM-based models, enables the model to capture contextual information from both directions, improving its ability to understand and represent the image data effectively.
  2. (2)
@@
    Selective Scan 2D : SS2D (Liu et al., 2024a) performs scanning operations in three directions: top to bottom, left to right, and in reverse direction. Each mamba block is placed to work independently within these directions. SS2D mirrors the self-attention process seen in transformers. It overcomes the limitations of bidirectional scan in ViM, but it also leads to a loss of patch continuity. To address this, SS2D incorporates a scan merge step, where representations from each scan direction are combined into a unified output.
  3. (3)
@@
    Continuous 2D Scan : Continuous 2D scan (Yang et al., 2024a) resolves the issue which is experienced in SS2D. It involves integrating direction-aware parameters into cross-scan mechanism and organizing patches accordingly. This approach ensures the preservation of patch continuity and maintains the contextual understanding of images. The continuous 2D scanning adds direction-aware parameter into data dependent parameter of SSM (B) which is expressed in Equation 13 and Equation 14.
    (13) $h^{\prime}_{k,i}=\overline{\textbf{{A}}}_{i}h_{k,i-1}+(\overline{\textbf{{B}}}_{i}+\overline{\mathbf{\Theta}}_{k,i})x_{i}$
    (14) $y^{\prime}_{i}=\mathbf{\sum_{k=1}^{4}}(\textbf{{C}}_{i}h^{\prime}_{k,i}+\textbf{{D}}x_{i}),\;\;\;\;\;\;y_{i}=y^{\prime}_{i}\mathbf{\odot}z_{i}$
  4. (4)
@@
    Zigzag Scan : The extension of continuous 2D Scan is zigzag scan (Hu et al., 2024) where the images are scanned with continuous scanning mechanism in both forward and backward direction. Zigzag scan is developed to enhance the continuity of patches in the images which are used for diffusion models such as ZigMa (Hu et al., 2024).
  5. (5)
@@
    Spatiotemporal Selective Scan : Spatiotemporal selective scan (Yang et al., 2024e) is used to scan on videos where the patches are unfolded on each frame along rows and columns and then concatenated with the frame sequence of $h_{i}^{t}\;\mathbf{\epsilon}\;\mathbb{R}^{C_{i}\times T(HW)}$ In this setup, scanning is done bidirectionally to know about temporal dependency. Parallelly, scanned patches are stacked around temporal axis to construct the spatial sequence in the form of $h_{i}^{s}\;\mathbf{\epsilon}\;\mathbb{R}^{C_{i}\times(HW)T}$, to integrate information of each pixel from all frames. In short, one scan focuses on scanning with time dependency along frames and the other focuses on scanning each pixel along the time axis.
  6. (6)
@@
    Local Scan : Local Scanning (Huang et al., 2024a) overcomes limitations of scanning methods in ViM and VMamba by preserving local dependencies in images through distinct local windows. This technique maintains the global context of the image without compromise. The authors suggest using $7\times 7$ and $2\times 2$ local windows to capture the local context while alternating the scan direction. Vertical and horizontal scans with direction flipping are used to grasp the global context of image tokens.
    Table 1. Summary of Scanning Mechanisms and Associated Models
@@
    BiDirectional Scan Vision Mamba (Zhu et al., 2024), MamMIL (Fang et al., 2024),
    Motion-Guided Dual-Camera Tracker (Zhang et al., 2024c)
    CAMS-Net (Khan et al., 2024)
    Selective Scan 2D(SS2D) VMamba (Liu et al., 2024a), MedMamba (Yue and Li, 2024), P-Mamba (Ye et al., 2024),
    Weak-Mamba-UNet (Wang and Ma, 2024), VM-UNET-V2 (Zhang et al., 2024d),
    LightM-UNet (Liao et al., 2024), HC-Mamba (Xu, 2024),
    H-vmunet (Wu et al., 2024a), Mamba-HUNet (Sanjid et al., 2024),
    UltraLight VM-UNet (Wu et al., 2024b), VM-UNet (Ruan and Xiang, 2024)
    VMambaMorph (Wang et al., 2024c), Mamba-UNet (Wang et al., 2024d),
    Swin-UMamba (Liu et al., 2024b), Semi-Mamba-UNet (Ma and Wang, 2024),
    MambaMIR (Huang et al., 2024b), VM-DDPM (Ju and Zhou, 2024)
    Spatiotemporal Selective Scan Vivim (Yang et al., 2024e)
    Zigzag Scan ZigMa (Hu et al., 2024)
    Local Scan LocalMamba (Huang et al., 2024a) , FreqMamba (Zhen et al., 2024)
    Efficient 2D Scan EfficientVMamba (Pei et al., 2024), FusionMamba (Peng et al., 2024)
    Continuous 2D Scan PlainMamba (Yang et al., 2024a)

### v3

@@ 3.1.2 Variants of U-Net
The integration of Mamba blocks into the U-Net [8] architecture has resulted in several variants which are designed to improve its performance. For instance, they can be added before the first encoder layer, after an encoder layer, within the skip connections or can even replace the whole encoder block in UNet architecture. Figure 13 presents a comparision of traditional U-Net and Mamba U-Net [78] architecture. Both networks share a standard UNet architecture having encoder and decoder blocks connected by skip connections but Mamba U-Net incorporates mamba blocks. Ruan et al. [115] proposed VM-UNet where images are converted into tokens using a patch embedding block. The encoder of the network contains two VSS blocks with patch merging block and skip connection in each layer of the encoder. The decoder contains a patch expanding block followed by a VSS block with skip connections added from the encoder. Finally, the representations from the decoder are then passed on to the final projection layer, which reconstructs the image back into its original size, i.e., the number of classes. Figure 14 illustrates the UNet-based architecture of VM-UNet.
Figure 14: Architecture of VM-UNet [115]
@@
Archit et al. [85] introduced ViM-UNet which uses ViM as its base Mamba. ViM-UNet features the ViM encoder with bidirectional SSM layers, while its decoder copies the design of UNEt TRansformers (UNETR) [129] using convolution and transposed convolution layers. To enhance efficiency, ViM-UNet introduces flexibility in its encoder sizes, such as tiny and small. ViM-UNet takes a different approach by excluding these skip connections and using ViM as the encoder. This key architectural change allows ViM-UNet to achieve efficient global feature extraction and improved segmentation performance, without the need for the skip connections that are important in the traditional UNet. Inspired by VM-UNet [115], Zhang et al. [116] proposed the second version called VM-UNet-V2. The model differs from VM-UNet by not using skip connections between subsequent encoder to decoder layers . Instead, features from each block of encoder are fused using the Semantics and Details Infusion block (SDI block). VSS captures contextual information within images, while the SDI module improves the integration of low-level and high-level features, leading to a comprehensive understanding of image features. By combining these elements, VM-UNetV2 maximizes the potential of SSMs within the UNet, thus offering a more efficient and powerful solution for segmentation tasks.
@@
Wang et al. [18] proposed LKM-UNet, a novel architecture designed for efficient 2D and 3D medical image segmentation. LKM-UNet uses the strengths of Mamba to achieve superior performance in both local and global modeling with linear time complexity. The use of large windows within SSM improves the receptive field compared to CNNs and Transformers. The network architecture includes a hierarchical and bidirectional Mamba block, enhancing spatial modeling capabilities by integrating Pixel-level (PiM) and Patch-level (PaM) SSMs. LKM-UNet excels in capturing both fine-grained details and long-range dependencies from the data. The model achieves efficient feature extraction and segmentation through a U-shaped network with encoder-decoder blocks, layers for downsampling and upsampling along with skip connections. Wu et al. [17] proposed High-order Vision Mamba UNet (H-vmunet) which enhances the 2D-selective-scan (SS2D) mechanism and also introduced higher-order interactions to reduce redundant information and improve extraction of local features. Traditional architecture like CNN and ViT face limitations in processing long sequences and capturing local features, respectively. The proposed architecture features a U-shaped structure with six layers, which incorporates High-order Visual State Space (H-VSS) module and it applies higher-order SS2D from layers three to six. Inclusion of Spatial Attention Bridge (SAB) and Channel Attention Bridge (CAB) modules in the architecture further enhances multilevel and multiscale information fusion, which is crucial for capturing detailed medical image features.
In Section 3.1.2, both Mamba-based and Transformer-based models aim to improve the U-Net architecture [8] by replacing standard convolutional layers. Mamba-based U-Nets incorporate blocks such as Vision Mamba (ViM)[97] and Visual State Space (VSS)[98], which improve the model’s ability to capture long-range dependencies efficiently. For example, the architecture of VM-UNet [115], shown in Figure 14, uses VSS blocks as core components within a U-Net-based structure. In contrast, Transformer-based U-Nets employ attention mechanisms, such as the Swin Transformer [130] and Vision Transformer (ViT)[126], to model global relationships in images. For instance, Swin U-Net[131] follows a similar architectural design to VM-UNet but replaces the VSS blocks with Swin Transformer blocks. While both approaches improve upon the original U-Net, Mamba-based models offer a better balance between accuracy and efficiency, achieving competitive segmentation performance with fewer parameters and reduced computational cost.
3.1.3 Hybrid Architectures
@@
CNNs often face challenges in capturing long-range dependencies due to their inherent focus on local features and computational complexity. SSMs have the ability to handle long sequences of data. The combination of Mamba with convolution plays a huge role in capturing local spatial information along with capturing long range dependencies from medical images. Some of the papers that demonstrate these hybrid approaches are as follows: Ma et al. [86] introduced U-Mamba, a hybrid CNN-SSM architecture integrating Mamba blocks within the encoder of U-Net, demonstrating superior performance over traditional CNN-based and transformer-based segmentation networks across various modalities and segmentation targets. Wang et al. [78] proposed Mamba-U-Net which integrates pure Vision mamba in the U-Net along with linear embedding and VSS block in the model. Xu et al. [83] developed HC-Mamba, which uses dilated convolution followed by depth-wise separable convolution along with Mamba. It improves the receptive field and reduces the parameters of the model. Another innovative model, SegMamba designed by Xing et al. [16], mimics a Transformer-based U-Net architecture, utilizing a series of TS-Mamba block and downsample block with residual connections inspired by He et al. [109]. TS-Mamba employs gated spatial convolution to analyze spatial relationships between 3D features, followed by the Tri-oriented Spatial Mamba (TOM) block, which replaces the traditional multi-headed attention layers in transformers. Gong et al. [20] proposed nnMamba, which features the Mamba-In-Convolution with Channel-Spatial Siamese (MICCSS) block. The backbone architecture of nnMamba leverages the MICCSS block to maintain computational efficiency. This helps in improving the model’s representational capacity and improving the performance on tasks that require a deep understanding of visual data.
Liu et al. [110] proposed Swin-UMamba which is a hybrid architecture that combines VSS with UNet for medical image segmentation tasks. Swin-UMamba integrates a Mamba-based encoder pretrained on ImageNet with a series of VSS blocks and upsample blocks. The architecture of Swin-UMamba includes patch embedding, VSS blocks, patch merging, up-sample blocks, residual blocks, and a 1x1 convolution for segmentation output which combines Mamba with traditional segmentation network components. By combining the strengths of Mamba-based models with segmentation network structures, Swin-UMamba achieves better results in medical image segmentation tasks, highlighting its effectiveness in improving pretrained models for improved segmentation accuracy.
@@
Wang et al. [79] proposed Weak-Mamba-UNet which uses a CNN-based UNet for extracting local features from data, a Swin Transformer-based SwinUNet for understanding global context and a VMamba based Mamba-UNet for modeling long-range dependencies. The results of Weak-Mamba-UNet shows that it solves inaccurate predictions caused by other approaches. Kirillov et al. proposed the Segment Anything Model (SAM) [132] which was originally done for 2D images. Inspired by this, Wang et al. [111] proposed Efficiently Adapting Segment Anything Model (SAM) which shows good performance on 3D medical images. The authors have modified the encoder of SAM by adding Tri-Plane Mamba (TP-Mamba) block on top of each ViT and the decoder of the model uses 3D convolution with instance normalization [133] and GELU activation function [134]. The authors have also used LoRA [135] to adapt weights of Multi-Head self Attention module. TP-Mamba shows better performance compared to transformer based models, U-Mamba [86] and other adapter-based algorithms for SAM such as SA-Med [136], MA-SAM [137].
Zhang et al. [112] proposed HMT-UNet which combines MambaVision Mixer with the transformer’s self attention block for segmentation of medical images. MambaVision mixer is a modification of the Mamba layer, where the first casual convolution is replaced with regular convolution and passed on to SSM. Initially, the encoder of UNet consists of four layers where the initial two layers uses two $3\times 3$ convolution with stride of two and downsampling uses batch normalized $3\times 3$ convolution with stride of two which scales down the resolution of the spatial dimension by half and doubles the channel. The layer 3 and layer 4 contains MambaVision Mixer and transformer’s self attention block. In the decoder block, the spatial dimensions are doubled and channel size is halved. The final stages of decoder use the same configuration as initial stages of encoder but instead of downsampling the image as in encoder, it linearly upsamples the image.
@@
VMRNN proposed by Tang et al. [113] is a recurrent cell that incorporates VSS block within an Long Short Term Memory (LSTM) [138] network. Mamba’s integration into spatial-temporal forecasting based on vision allows robust sequential modeling. The paper introduces two novel architectures: VMRNN-B and VMRNN-D. These architectures excel at extracting spatiotemporal features, establishing a new strong foundation for spatiotemporal forecasting. Spatiotemporal predictive learning differs from traditional image-level vision tasks by predicting future video frames based on past sequences. VMRNN tackles this challenge by processing individual frames, segmenting them into smaller patches. These patches get flattened before feeding them into a patch embedding layer for initial processing. The VMRNN layer then leverages these transformed patches alongside past states to capture the crucial spatiotemporal features necessary for predicting the next frame. The hybrid architecture of VSS and LSTM is attributed to the model’s capability to learn and leverage global spatial dependencies with linear complexity, enabling a more refined understanding of spatiotemporal dynamics.
@@
Ding et al. [139] combines Mamba with Graph Neural Network (GNN) to capture global and local tissue spatial relationships respectively in Whole Slide Images (WSI). The model leverages a message-passing GNN, specifically a Graph Attention Network (GAT) proposed by Velickovic et al. [140], to process a hierarchical graph constructed from both cell-level and tile-level graphs. Graph convolution operation is performed using GAT which assigns importance scores to neighboring nodes, allowing the model to focus on the most relevant information for the current node during the convolution process.
@@
Nasiri et al. [22] proposed Vim4Path which used ViM architecture within the DINO framework proposed by Caron et al. [141], for learning representations in computational pathology. ViM is modified to accept arbitrary input image sizes using positional embedding interpolation, making it adaptable within DINO for Self-Supervised Learning (SSL). The study benchmarks the Camelyon16 dataset [63], extracting image patches from Whole Slide Images (WSIs) without labels for training the ViM encoder using the DINO framework. Yang et al. [114] introduced the MambaMIL framework, which combines Mamba with Multiple Instance Learning (MIL) to enhance long sequence modeling in computational pathology. The core component of the MambaMIL framework is Sequence Reordering Mamba (SR-Mamba), which is designed to be aware of the order and distribution of instances within long sequences. The MambaMIL framework partitions tissue regions into a sequence of patches, maps these patches into instance features, reduces feature dimension through linear projection, utilizes stacked SR-Mamba modules for handling long sequences, and finally employs an aggregation module to obtain bag-level representations for downstream tasks.
This Section 3.1.3 has explored various architectures that combine Mamba with CNNs, Transformers, GNNs, and other models to tackle complex tasks like medical imaging. Relying on a single architecture often falls short in effectively capturing both fine-grained local features and long-range dependencies, making the combination of complementary models increasingly important. Mamba offers efficient sequence modeling with low computational cost, making it a valuable addition to these architectures. In CNN-based models such as U-Mamba [86] and HC-Mamba [83], it enhances spatial feature learning and long-range dependency capture. In Transformer-based models such as TP-Mamba [111] and HMT-UNet [112], Mamba reduces the computational burden of self-attention while maintaining global context understanding. In recurrent models like VMRNN [113], it strengthens temporal modeling for video forecasting, while in GNNs, it improves hierarchical spatial reasoning through attention-driven message passing. Common components in these combinations such as attention modules, gating mechanisms, residual connections, patch embeddings, and lightweight convolutions, demonstrate a clear trend toward leveraging Mamba’s strengths to boost performance and efficiency across diverse deep learning tasks.
3.2 Scanning
(a) Illustrations in LABEL:fig:scanninga, LABEL:fig:scanningb, LABEL:fig:scanningc & LABEL:fig:scanningd present Different Scanning Mechanisms used in Mamba-based Architectures
@@ 3.2 Scanning
Attention mechanisms, especially self-attention have a quadratic time complexity causing computational costs to grow quadratically with sequence length. In contrast, scanning operations generally have linear time complexity, making them more efficient for long sequences. The scan operation involves calculating an array, like the prefix sum, where each value is determined by using the previously calculated value and the current input. Similarly, the recurrent form of SSM can be viewed as a scan operation. Scanning is a crucial component in mamba, especially when handling multidimensional inputs. The selection of the scanning mechanism in Mamba models is crucial as it enhances efficiency and provides important information. 16(a) provides visualization of various scanning mechanisms employed in mamba-based architectures. This visualization highlights the diversity of scanning approaches integrated into mamba models. Table 1 summarizes the various scanning mechanisms and the models associated with each mechanism. The scanning methodologies used in mamba models are detailed as follows:
  1. 1.
@@
    Bidirectional Scan : In bidirectional scan (forward and backward scan) [97], after tokenizing image patches, they are processed through the forward SSM. Simultaneously, the same tokenized representations of images are independently processed through the backward SSM. This scanning mechanism primarily used in ViM-based models, enables the model to capture contextual information from both directions, improving its ability to understand and represent the image data effectively.
  2. 2.
@@
    Selective Scan 2D : SS2D [98] performs scanning operations in three directions: top to bottom, left to right, and in reverse direction. Each mamba block is placed to work independently within these directions. SS2D mirrors the self-attention process seen in transformers. It overcomes the limitations of bidirectional scan in ViM, but it also leads to a loss of patch continuity. To address this, SS2D incorporates a scan merge step, where representations from each scan direction are combined into a unified output.
  3. 3.
@@
    Continuous 2D Scan : Continuous 2D scan [102] resolves the issue which is experienced in SS2D. It involves integrating direction-aware parameters into cross-scan mechanism and organizing patches accordingly. This approach ensures the preservation of patch continuity and maintains the contextual understanding of images. The continuous 2D scanning adds direction-aware parameter into data dependent parameter of SSM (B) which is expressed in Equation 13 and Equation 14.
    $h^{\prime}_{k,i}=\overline{\textbf{{A}}}_{i}h_{k,i-1}+(\overline{\textbf{{B}}}_{i}+\overline{\mathbf{\Theta}}_{k,i})x_{i}$ (13)
    $y^{\prime}_{i}=\mathbf{\sum_{k=1}^{4}}(\textbf{{C}}_{i}h^{\prime}_{k,i}+\textbf{{D}}x_{i}),\;\;\;\;\;\;y_{i}=y^{\prime}_{i}\mathbf{\odot}z_{i}$ (14)
  4. 4.
@@
    Zigzag Scan : The extension of continuous 2D Scan is zigzag scan [99] where the images are scanned with continuous scanning mechanism in both forward and backward direction. Zigzag scan is developed to enhance the continuity of patches in the images which are used for diffusion models such as ZigMa [99].
  5. 5.
@@
    Spatiotemporal Selective Scan : Spatiotemporal selective scan [73] is used to scan on videos where the patches are unfolded on each frame along rows and columns and then concatenated with the frame sequence of $h_{i}^{t}\;\mathbf{\epsilon}\;\mathbb{R}^{C_{i}\times T(HW)}$ . In this setup, scanning is done bidirectionally to know about temporal dependency. Parallelly, scanned patches are stacked around temporal axis to construct the spatial sequence in the form of $h_{i}^{s}\;\mathbf{\epsilon}\;\mathbb{R}^{C_{i}\times(HW)T}$, to integrate information of each pixel from all frames. In short, one scan focuses on scanning with time dependency along frames and the other focuses on scanning each pixel along the time axis.
  6. 6.
@@
    Local Scan : Local Scanning [100] overcomes limitations of scanning methods in ViM and VMamba by preserving local dependencies in images through distinct local windows. This technique maintains the global context of the image without compromise. The authors suggest using $7\times 7$ and $2\times 2$ local windows to capture the local context while alternating the scan direction. Vertical and horizontal scans with direction flipping are used to grasp the global context of image tokens.
    Table 1: Summary of Scanning Mechanisms and Associated Models
@@
    BiDirectional Scan Vision Mamba [97], MamMIL [142],
    Motion-Guided Dual-Camera Tracker [143]
    CAMS-Net [89]
    Selective Scan 2D(SS2D) VMamba [98], MedMamba [21], P-Mamba [71],
    Weak-Mamba-UNet [79], VM-UNET-V2 [116],
    LightM-UNet [80], HC-Mamba [83],
    H-vmunet [17], Mamba-HUNet [87],
    UltraLight VM-UNet [81], VM-UNet [115]
    VMambaMorph [75], Mamba-UNet [78],
    Swin-UMamba [110], Semi-Mamba-UNet [84],
    MambaMIR [23], VM-DDPM [74]
    Spatiotemporal Selective Scan Vivim [73]
    Zigzag Scan ZigMa [99]
    Local Scan LocalMamba [100] , FreqMamba [144]
    Efficient 2D Scan EfficientVMamba [101], FusionMamba [145]
    Continuous 2D Scan PlainMamba [102]